### v2

# Unraveling time-varying causal effects of multiple exposures: integrating Functional Data Analysis with Multivariable Mendelian Randomization

Abstract
Mendelian Randomization is a widely used instrumental variable method for assessing causal effects of lifelong exposures on health outcomes. Many exposures, however, have causal effects that vary across the life course and often influence outcomes jointly with other exposures or indirectly through mediating pathways. Existing approaches to multivariable Mendelian Randomization assume constant effects over time and therefore fail to capture these dynamic relationships. We introduce Multivariable Functional Mendelian Randomization (MV-FMR), a new framework that extends functional Mendelian Randomization to simultaneously model multiple time-varying exposures. The method combines functional principal component analysis with a data-driven cross-validation strategy for basis selection and accounts for overlapping instruments and mediation effects. Through extensive simulations, we assessed MV-FMR’s ability to recover time-varying causal effects under a range of data-generating scenarios and compared the performance of joint versus separate exposure effect estimation strategies. Across scenarios involving nonlinear effects, horizontal pleiotropy, mediation, and sparse data, MV-FMR consistently recovered the true causal functions and outperformed univariable approaches. To demonstrate its practical value, we applied MV-FMR to UK Biobank data to investigate the time-varying causal effects of systolic blood pressure and body mass index on coronary artery disease. MV-FMR provides a flexible and interpretable framework for disentangling complex time-dependent causal processes and offers new opportunities for identifying life-course critical periods and actionable drivers relevant to disease prevention.
Keywords Mendelian randomization $\cdot$ functional data analysis $\cdot$ causal inference $\cdot$ longitudinal data $\cdot$ life-course epidemiology
1 Introduction
Establishing causal relationships between exposures and health outcomes remains a central challenge in epidemiological research. Mendelian Randomization (MR) is a widely used method for causal inference that leverages genetic variants as instrumental variables (IVs) to estimate the causal effect of an exposure on an outcome [sanderson2022mendelian, Lawlor2008]. Traditional MR methods generally assume that a single exposure exerts a constant causal effect. Since genetic variants are fixed at conception, estimates from traditional MR are typically interpreted as reflecting the average effect over a lifetime of genetically predicted exposure across the entire life course, reflecting cumulative rather than age-specific or period-specific influences [burgess2021interpretation]. However, many exposures exhibit time-varying effects, meaning that their causal impact on outcomes can change across different stages of life [Kaneko2023]. For example, the effect of body mass index on cardiovascular disease risk may differ between early adulthood and older age. This variation can occur due to evolving physiological processes or age-specific biological mechanisms [Palmer2022, Karlsson2020]. To capture these dynamics, analytical methods must move beyond single time-point estimates and instead model how effects evolve over time. This approach allows for the identification of age-specific or life course-specific influences that may be overlooked when relying on a single measurement of exposures [wagner2021time, shi2022mr_timevarying]. Moreover, emerging evidence also indicates that genetic associations with certain exposures can vary at different time points in the life course (within a population), meaning that MR estimates may reflect the effect of exposure at particular life periods rather than uniformly over time [sanderson2022timevarying, jiang2021age]. Accounting for these time-varying genetic effects allows researchers to identify critical periods when an exposure has the greatest impact on health outcomes [Power2024Mendelian]. Incorporating these dynamics into MR analyses, therefore, enhances causal inference for time-dependent exposures and can inform the design and timing of preventive interventions. In response to this challenge, recent work has focused on estimating time-varying causal effects using MR. Most existing methods discretize time, treating exposures measured at fixed points as separate variables, ignoring the continuous nature of exposure trajectories. Tian et al. [Tian2024] proposed a continuous-time MR framework, called Univariable Functional Mendelian Randomization (U-FMR), which models the exposure effect as a smooth function of time using functional data analysis [ramsay2005fda]. Their method uses Functional Principal Component Analysis (FPCA) to reduce the dimensionality of longitudinal exposure data and integrates these low-dimensional representations into a MR model. This framework provides a flexible and theoretically robust approach for inferring how causal effects evolve over the life course, given longitudinal data.
Despite these advances, existing MR methods that investigate the life course effects of exposure have focused exclusively on single exposures [Tian2024, cao2016timevarying, richardson2020earlylater, shi2021timevarying, sanderson2022timevarying, richardson2023timevarying]. However, in real-world scenarios, multiple exposures often act simultaneously or influence each other over the life course, particularly when they are biologically related. Multivariable MR (MV-MR) extends the univariable MR (U-MR) framework to jointly estimate the causal effects of multiple exposures using a set of genetic variants as instruments [sanderson2019examination]. Unlike U-MR, which captures only the total effect of a single exposure, MV-MR can disentangle the direct effects of many exposures, provided that the genetic instruments are sufficiently associated with each exposure throughout the life course. However, existing MV-MR methods assume that causal effects are constant over time.
To address the limitation of time-invariant effects, we propose Multivariable Functional Mendelian Randomization (MV-FMR), which extends U-FMR [Tian2024] to simultaneously model multiple time-varying exposures. Our key methodological contributions include: (1) a data-driven cross-validation procedure for selecting the number of basis components for each exposure, (2) joint modeling of time-varying exposures that share common genetic pathways, (3) handling of mediation pathways between time-varying exposures, and (4) implementation for both continuous and binary outcomes. To evaluate the performance of MV-FMR, we conducted extensive simulation studies under three main scenarios: horizontal pleiotropy through overlapping genetic instruments, mediation pathways where one exposure influences another, and null effects to assess specificity. We further performed sensitivity analyses examining different functional forms for the exposure-outcome effects (linear and nonlinear effects) and outcome types (continuous and binary). Results demonstrated that MV-FMR accurately recovered time-varying causal effects, consistently outperforming univariable functional MR approaches. We apply MV-FMR to UK Biobank data to estimate the time-varying causal effects of systolic blood pressure and body mass index on coronary artery disease risk across adulthood, demonstrating how the method can identify critical age periods and disentangle direct effects from mediation pathways that would be missed by traditional approaches.
The remainder of the paper is organized as follows. In Section 2, we review the background on MV-MR and U-FMR, and introduce our proposed extension, including its statistical formulation and estimation strategy. Section 3 presents simulation studies evaluating the performance of MV-FMR compared to the univariable approach. In Section 4, we apply the method in a proof-of-concept application using UK Biobank data. Finally, Section 5 concludes with a discussion of the results, limitations, and directions for future research.
2 Methods
2.1 Background on Multivariable Mendelian Randomization (MV-MR)
Multivariable MR (MV-MR) generalizes the classical MR framework to jointly estimate the direct causal effects of multiple exposures on an outcome using genetic variants as instrumental variables [sanderson2021multivariable]. While univariable MR captures the total effect of a single exposure, MV-MR disentangles the direct effect of each exposure while accounting for the influence of the others. Estimating multiple exposures simultaneously is often desirable, for example, when exposures are biologically or clinically related, or when one exposure may mediate the effect of another on the outcome. By modeling multiple exposures simultaneously, MV-MR can disentangle overlapping causal mechanisms and improve robustness to pleiotropy, provided that pleiotropic effects operate only through the modeled exposures. The validity of MV-MR requires an extension of the standard IV assumptions to the multivariable setting:
  1. 1.
    Relevance (MV-IV1): each instrument must be strongly associated with at least one exposure, conditional on the others.
  2. 2.
    Exclusion Restriction (MV-IV2): instruments must affect the outcome only through the included exposures.
  3. 3.
    Independence (MV-IV3): each instrument must be independent of any confounders of the exposures-outcome relationship.
2.2 Overview of the Univariate Functional Mendelian Randomization (U-FMR) with continuous outcome
We begin by briefly showing the U-FMR model proposed by [Tian2024] as a foundation for our multivariable extension presented in 2.3. The setting involves a continuous outcome $Y$, a time-varying exposure $X(t)$, and a set $P$ of genetic instruments $\bm{G}$. Inference on time-varying causal effects relies on life-course variation in the associations between the genetic instruments and the exposure. When such associations evolve over time, they provide information that enables the identification of how the causal effect of the exposure on the outcome changes across the life course. Under this setting, the instruments act as valid proxies for the exposure across the observation period, enabling the estimation of the time-varying causal effect to be estimated. The causal effect is therefore modeled as a smooth function $\beta(t)$, capturing how the impact of $X(t)$ on $Y$ evolves over the observed time interval. This leads to the structural model:
$Y=\beta_{0}+\int_{0}^{T}\beta(t)X(t)\,dt+\varepsilon.$ (1)
Because exposure measurements are often sparse and noisy, individual trajectories $X(t)$ are treated as functional objects and modeled in a linear additive form using a basis expansion derived from Functional Principal Component Analysis (FPCA):
$X(t)\approx\mu(t)+\sum_{k=1}^{K}\xi_{k}\phi_{k}(t),$ (2)
with $K$ denoting the number of selected principal components, $\mu(t)$ is the mean exposure trajectory, $\phi_{k}(t)$ are eigenfunctions of the covariance operator, and $\xi_{k}$ are subject-specific FPCA scores. Similarly, the causal effect function $\beta(t)$ is represented using as basis functions the eigenfunctions of the FPCA:
$\beta(t)=\sum_{k=1}^{K}\alpha_{k}\phi_{k}(t),$ (3)
where $\alpha_{k}$ are the corresponding coefficients to be estimated. Substituting the expansions for $X(t)$ and $\beta(t)$ from equations (2) and (3) into the structural model (1), the equation simplifies to:
$Y\approx\beta_{0}^{*}+\sum_{k=1}^{K}\beta^{*}_{k}\xi_{k},$ (4)
where $\beta_{0}^{*}=\beta_{0}+\int_{0}^{T}\beta(t)\mu(t)dt$ and each pseudo-exposure coefficient $\beta^{*}_{k}$ is defined as:
$\beta^{*}_{k}=\int_{0}^{T}\beta(t)\phi_{k}(t)\,dt.$ (5)
This formulation reduces the original functional model to a standard linear regression on the subject-specific FPCA scores $\xi_{k}$, making it more tractable for estimation and inference in settings with sparse or noisy exposure measurements.
The coefficients $\beta^{*}_{k}$ are treated as pseudo-exposures, and estimated using the continuously updating Generalized Method of Moments (GMM) [hansen1996finite], where genetic variants $\bm{G}$ are used as instruments for the FPCA scores $\xi_{k}$. GMM is particularly well-suited for this setting as it performs robustly under weak instrument conditions commonly encountered with genetic instrumental variables. All variables are mean-centered prior to estimation to ensure numerical stability. The sample moment function is defined as:
$\mathbf{g}(\boldsymbol{\beta}^{*})=\frac{1}{n}\mathbf{G}^{T}(\mathbf{Y}-\boldsymbol{\xi}\boldsymbol{\beta}^{*}),$ (6)
and the GMM estimator is obtained by minimizing the continuously updated objective function:
$\boldsymbol{\hat{\beta}}^{*}=\arg\min_{\boldsymbol{\beta}^{*}}\mathbf{g}(\boldsymbol{\beta}^{*})^{T}\mathbf{W}(\boldsymbol{\beta}^{*})^{-1}\mathbf{g}(\boldsymbol{\beta}^{*}),$ (7)
where the weighting matrix is $\mathbf{W}(\boldsymbol{\beta}^{*})=\frac{1}{N}\mathbf{G}^{T}\text{diag}(\mathbf{Y}-\boldsymbol{\xi}\boldsymbol{\beta}^{*})^{2}\mathbf{G}$. Under standard regularity conditions, $\boldsymbol{\hat{\beta}}^{*}$ is consistent and asymptotically normally distributed.
Once the coefficients $\beta^{*}_{k}$ are estimated, the functional effect $\beta(t)$ can be reconstructed by inverting the relationship between $\beta(t)$ and the FPCA basis using:
$\beta(t)\approx\sum_{k=1}^{K}\beta^{*}_{k}\phi_{k}(t).$ (8)
This framework provides the basis for extending the model to multiple exposures, as illustrated in Figure 1. While U-FMR (panel a) estimates total effects separately, multivariable functional MR (panel b) jointly models both exposures to recover direct causal effects.
Figure 1: Comparison of (a) Univariable functional MR (independent effects of $X_{1}$ and $X_{2}$) with (b) Multivariable functional MR (joint effects of $X_{1}$ and $X_{2}$). Dashed arrows represent confounding pathways.
2.3 Multivariable Functional Mendelian Randomization (MV-FMR) with continuous outcome
Building on the U-FMR framework described in 2.2 and the causal structure illustrated in Figure 1, we extend the model to a multivariate setting to accommodate multiple time-varying exposures, each with its own dynamic causal effect. Consider two exposures, $X_{1}(t)$ and $X_{2}(t)$, observed over $t\in[0,T]$. While we focus on two exposures for simplicity, the proposed framework naturally generalizes to higher dimensions. We assume that each exposure trajectory can be represented using FPCA:
$X_{j}(t)=\mu_{j}(t)+\sum_{k=1}^{K_{j}}\xi_{j,k}\,\phi_{j,k}(t)+\epsilon_{j}(t),\quad j\in\{1,2\},$ (9)
where $\mu_{j}(t)$ is the mean trajectory, $\phi_{j,k}(t)$ are eigenfunctions and $\xi_{j,k}$ are FPCA scores, for exposure $j$. The outcome is modeled as a function of both exposures:
$Y=\beta_{0}+\int_{0}^{T}\beta_{1}(t)X_{1}(t)\,dt+\int_{0}^{T}\beta_{2}(t)X_{2}(t)\,dt+g_{Y}(U,\epsilon_{Y}),$ (10)
where $\beta_{j}(t)$ denotes the direct time-varying causal effect of exposure $j$, and $g_{Y}(U,\epsilon_{Y})$ captures unmeasured confounding and random noise.
By substituting the FPCA expansion (9) into (10), we can rewrite the model as:
$Y$ $=\beta_{0}+\int_{0}^{T}\beta_{1}(t)\left[\mu_{1}(t)+\sum_{k_{1}=1}^{K_{1}}\xi_{1,k_{1}}\,\phi_{1,k_{1}}(t)+\epsilon_{1}(t)\right]dt+\int_{0}^{T}\beta_{2}(t)\left[\mu_{2}(t)+\sum_{k_{2}=1}^{K_{2}}\xi_{2,k_{2}}\,\phi_{2,k_{2}}(t)+\epsilon_{2}(t)\right]dt+g_{Y}(U,\epsilon_{Y})$ (11)
$=\underbrace{\beta_{0}+\int_{0}^{T}\left[\beta_{1}(t)\mu_{1}(t)+\beta_{2}(t)\mu_{2}(t)\right]dt}_{\beta_{0}^{*}}+\sum_{k_{1}=1}^{K_{1}}\underbrace{\left[\int_{0}^{T}\beta_{1}(t)\phi_{1,k_{1}}(t)\,dt\right]}_{\beta_{k_{1}}^{*}}\xi_{1,k_{1}}+\sum_{k_{2}=1}^{K_{2}}\underbrace{\left[\int_{0}^{T}\beta_{2}(t)\phi_{2,k_{2}}(t)\,dt\right]}_{\beta_{k_{2}}^{*}}\xi_{i2,k_{2}}+\tilde{\epsilon}$,
where $\tilde{\epsilon}$ collects residual terms, including measurement noise and unobserved confounding. This formulation reduces the multivariate functional problem to an MV-MR, in which the FPCA scores $\xi_{j,k}$ act as pseudo-exposures, and the corresponding coefficients $\beta_{j,k}^{*}$ can be estimated using genetic instruments.
2.3.1 Estimation of component effects
Each exposure trajectory is represented using FPCA, with scores estimated via the PACE method [Yao2005] to accommodate sparse and irregular measurements. Full mathematical details of the FPCA representation and PACE estimation are provided in Supplementary Method 1. Let $\mathbf{G}=[\mathbf{G}_{1},\mathbf{G}_{2},\mathbf{G}_{12}]\in\mathbb{R}^{N\times P}$ denote the stacked matrix of genetic instruments, where: $\mathbf{G}_{j}\in\mathbb{R}^{N\times P_{j}}$ are instruments specific to $X_{j}(t)$, and $\mathbf{G}_{12}\in\mathbb{R}^{N\times P_{12}}$ are pleiotropic instruments influencing both $X_{1}(t)$ and $X_{2}(t)$. The estimation of $\boldsymbol{\beta}^{*}_{1}$ and $\boldsymbol{\beta}^{*}_{2}$ follows the GMM framework described in Section 2.2, extending it to jointly model both exposures. The stacked FPCA scores $\boldsymbol{\xi}=[\boldsymbol{\xi}_{1},\boldsymbol{\xi}_{2}]\in\mathbb{R}^{N\times(K_{1}+K_{2})}$ and coefficient vector $\boldsymbol{\beta}^{*}=[\boldsymbol{\beta}^{*}_{1},\boldsymbol{\beta}^{*}_{2}]$ are used in the GMM objective function with $\mathbf{G}$ as the full instrument matrix. This joint estimation allows the model to account for correlation between exposures through overlapping instruments while maintaining the asymptotic properties established for the univariable case.
2.3.2 Data-driven selection of FPCA components
To approximate each exposure trajectory, the FPCA expansion is truncated at $K_{j}$ components. The number of retained FPCs $K_{j}$ critically determines the flexibility of the causal effect estimation. We select $K_{j}$ using a data-driven k-fold cross-validation procedure that optimizes predictive performance. This approach in this setting outperforms variance-based criteria, which often retain too few components and fail to capture higher-order effects. Model performance is evaluated via mean squared error (MSE). This approach balances model complexity and generalizability, capturing meaningful functional effects while minimizing overfitting.
2.3.3 Reconstructing the coefficient functions
The time-varying causal effect functions $\beta_{j}(t)$ are reconstructed from the estimated pseudo-exposure coefficients using the eigenfunction basis. Let $\beta_{k_{j}}^{*}=\int_{0}^{T}\beta_{j}(t)\phi_{j,k}(t)dt\,$ then the reconstruction formula is given by:
$\hat{\beta}_{j}(t)=\boldsymbol{\phi}^{T}_{j}(t)\boldsymbol{\hat{\beta}}^{*}_{j},$ (12)
where $\boldsymbol{\phi}_{j}(t)=[\phi_{j,1}(t),\ldots,\phi_{j,K_{j}}(t)]^{T}$ denotes the vector of eigenfunctions and $\boldsymbol{\hat{\beta}}^{*}_{j}\in\mathbb{R}^{K_{j}}$ is the corresponding vector of estimated coefficients.
2.4 Multivariable Functional Mendelian Randomization (MV-FMR) with binary outcome
We also propose an extension of the MV-FMR framework to binary outcomes by modifying the structural equation and estimation procedure while preserving the functional representation of exposures. The outcome model in 10 becomes:
$\text{logit}(\mathbb{P}(Y=1))=\beta_{0}+\int_{0}^{T}\beta_{1}(t)X_{1}(t)\,dt+\int_{0}^{T}\beta_{2}(t)X_{2}(t)\,dt+g_{Y}(U).$ (13)
The FPCA representation of exposures in 9 and the reduction to pseudo-exposures in 11 remain unchanged, yielding a logistic model for the FPCA scores.
For binary outcomes, we apply a two-stage residual inclusion (2SRI) control function approach to account for endogeneity in the FPCA scores. This method is preferred for nonlinear models over standard two-stage predictor substitution (2SPS) approaches, which replace endogenous variables with their predicted values. In the context of nonlinear models such as logistic regression, the 2SPS estimator produces inconsistent estimates because the nonlinear link function invalidates the substitution of predicted values, thereby distorting the conditional mean structure [terza2008two, wooldridge2015control]. In contrast, the 2SRI approach controls for endogeneity by including first-stage residuals as additional regressors. Following the notation introduced in 2.3, in the first stage, FPCA scores are regressed on instruments:
$\boldsymbol{\xi}=\mathbf{G}\boldsymbol{\Gamma}+\mathbf{V},\quad\hat{\mathbf{V}}=\boldsymbol{\xi}-\mathbf{G}\hat{\boldsymbol{\Gamma}}$, (14)
where $\hat{\mathbf{V}}$ are the residuals capturing variation in FPCA scores not explained by the instruments. The second stage includes both scores and residuals in a logistic regression:
$\text{logit}(\mathbb{P}(Y=1))=\mathbf{1}_{n}\boldsymbol{\beta}_{0}+\boldsymbol{\xi}\boldsymbol{\beta}^{*}+\hat{\mathbf{V}}\boldsymbol{\alpha}$ (15)
where $\boldsymbol{\hat{\beta}^{*}}$ estimates the direct effect of the FPCA scores on the outcome and $\boldsymbol{\alpha}$ captures the correlation between the FPCA scores and unobserved confounders, correcting for endogeneity. Robust standard errors are used to account for the estimation of $\hat{\mathbf{V}}$ in the first stage. Under standard regularity conditions and correct specification of the first-stage model, $\boldsymbol{\hat{\beta}^{*}}$ is consistent and asymptotically normal.
The data-driven component selection procedure (2.3.2) and the reconstruction of $\hat{\beta}_{j}(t)$ (2.3.3) follow the same approach as for continuous outcomes, with the exception that the area under the ROC curve (AUC) replaces MSE as the performance metric for cross-validation in the component selection procedure.
3 Simulation Studies
We conduct comprehensive simulation studies to evaluate the performance of our MV-FMR approach. The simulations assess the method’s ability to recover time-varying causal effects under various data-generating scenarios and compare the performance of joint versus separate exposure effects estimation strategies.
3.1 Simulation scenarios
We designed three distinct simulation scenarios (see Figure 2), including both exposure-specific and overlapping instruments (pleiotropic), to assess the method’s ability to accurately estimate direct causal effects.
  • Scenario 1: Pleiotropy model (Figure 2, left panel): Exposures have an independent pleiotropic pathway from instruments to the outcome. This scenario evaluates the ability of MV-FMR to disentangle direct causal effects in the presence of horizontal pleiotropy.
  • Scenario 2: Null effect control (Figure 2, center left): Same genetic architecture as in Scenario 1, but the second exposure does not have a causal effect on the outcome $(\beta_{2}(t)=0\forall t)$. This scenario evaluates the type I error, ensuring that MV-FMR does not detect spurious causal effects.
  • Scenario 3: Mediation model (Figure 2, right left): $X_{1}(t)$ causally influences $X_{2}(t)$, and both affect the outcome $Y$. This scenario evaluates the ability of MV-FMR to recover direct effects while accounting for mediation.
Figure 2: Illustration of the three simulation scenarios used to evaluate MV-FMR. Scenario 1: Pleiotropy model (left). Scenario 2: Null effect control (center). Scenario 3: Mediation model (right).
These three scenarios reflect common epidemiological challenges that arise when investigating multiple correlated risk factors [sanderson2022timevarying]. Horizontal pleiotropy is frequently observed when genetic variants influence multiple physiological pathways, potentially violating the exclusion restriction in univariable MR analysis. Mediation pathways are pervasive in chronic disease epidemiology, where one risk factor operates both directly on the outcome and indirectly through downstream exposures. In such cases, U-FMR estimates the total causal effect, whereas MV-FMR recovers the direct effect by conditioning on the mediator. Finally, it is important to assess specificity in settings where some exposures have true null effects. In practice, analyses of joint time-varying exposures may include variables that are correlated with causal risk factors but do not themselves affect disease risk, and we aim to verify that the model does not introduce spurious non-null estimates. Correctly distinguishing these relationships is essential for identifying modifiable targets for intervention and avoiding spurious causal inferences.
3.2 Data generating process
We simulate individual-level functional trajectories $X_{j}(t)$ for each exposure $j=1,2$ over the time domain $t\in[0,50]$, according to the following exposure model:
$X_{j}(t)=\sum_{l=1}^{P}\alpha_{l}(t)G_{l}+[U_{0j}+U_{j}(t)]+\varepsilon_{j}(t),$ (16)
where $G_{l}\sim\text{Binomial}(2,0.3)$ denotes the $l$-th genetic variant. The time-varying genetic effect $\alpha_{l}(t)$ determines the influence of each variant on exposure $j$ over the observation period, while $U_{0j}$ and $U_{j}(t)$ denote time-fixed and time-varying confounders affecting the exposure. The residual term $\varepsilon_{j}(t)$ accounts for exposure-specific noise and measurement error. Both $U_{j}(t)$ and $\varepsilon_{j}(t)$ are modeled as discretized Wiener processes, and $U_{0j}\sim\mathcal{N}(0,1)$. To emulate sparse longitudinal data, for each exposure, each individual is observed at $nSparse=10$ randomly selected time points over $[0,50]$. The time-varying genetics effect on exposures $\alpha_{l}(t)$ are generated:
$\alpha_{l}(t)=a_{l}+b_{l}\cdot t,\quad a_{l}\sim\text{Uniform}(-0.1,0.1),\quad b_{l}\sim\text{Uniform}(-0.004,0.004).$ (17)
Only for Scenario 3, $X_{1}(t)$ influences $X_{2}(t)$ and then is generated as:
$X_{2}(t)=\sum_{l=1}^{P}\alpha_{l}(t)G_{l}+\gamma\cdot X_{1}(t)+[U_{0j}+U_{j}(t)]+\varepsilon_{X,2}(t),$ (18)
where $\gamma$ controls the strength of the mediation pathway. In the baseline simulation $\gamma=0.3$, then additional simulations were performed varying $\gamma$ to assess its impact on MV-FMR performance.
The outcome $Y_{i}$ is simulated according to:
$Y_{i}=\beta_{0}+\int_{0}^{T}\beta_{1}(t)X_{i1}(t)dt+\int_{0}^{T}\beta_{2}(t)X_{i2}(t)dt+\varepsilon_{Y_{i}},\varepsilon_{Y_{i}}\sim\mathcal{N}(0,1),$ (19)
with integration approximated numerically over a regular grid. Three functional forms for the exposure-outcome effects $\beta_{j}(t)$ are considered:
  • Null effect: $\beta_{j}(t)=0$,
  • Linear effect: $\beta_{j}(t)=0.02\cdot t$,
  • Quadratic effect: $\beta_{j}(t)=0.002\cdot t^{2}-0.11\cdot t+0.5$.
Functional principal components are calculated using the fdapace package [fdapace]. We compare the performance of MV-FMR against U-FMR applied separately to each exposure. Performance is assessed using the Integrated Squared Error (ISE), which quantifies the average squared deviation between estimated and true effect functions, and the pointwise 95% coverage rate, defined as the proportion of time points where the 95% confidence intervals contain the true effect. Each simulation scenario is replicated $R=500$ times with a sample size $N=5000$. The number of instruments is set to $P_{1}=P_{2}=100-P12$, where $P_{12}=p_{12}\cdot 100$ with $p_{12}=0.15$ in the baseline simulations. Additional simulations vary the degree of instrument overlap across exposures, considering values of $p_{12}$ equal to 0.15, 0.3, 0.5 and 0.8, to explore how different levels of overlap affect MV-FMR performance. Instrument strength is measured using the conditional F-statistic for each principal component [Sanderson2021-Fstat].
3.3 Simulation results
We first compared the performance of MV-FMR against U-FMR across the three scenarios (Figure 2), assuming that both exposure-outcome causal effects are linear. The ISE results for models with continuous and binary outcomes are reported in Table 1, with coverage probabilities and conditional F-statistics provided in the Supplementary Tables 1 and 2. MV-FMR demonstrated superior performance across all evaluated scenarios, consistently achieving a lower mean ISE compared to the univariate approach. Specifically, in Scenario 1, MV-FMR showed robustness to overlapping variants and in Scenario 2 correctly identified the null association effect of $X_{2}$ on the outcome, while the univariable method yielded false positive findings. Finally, in Scenario 3, MV-FMR successfully disentangled direct causal effects from mediated effects for both exposures, whereas the univariable approach could not distinguish between these pathways, resulting in increased errors. When examining performance by outcome type, MV-FMR maintained its superior accuracy across both continuous and binary endpoints. Although ISE values were greater for binary outcomes, reflecting the inherent increased variability in dichotomous responses compared to continuous measures, MV-FMR consistently outperformed the univariable method regardless of outcome specification.
Continuous Outcome Binary Outcome
Scenario MV-FMR U-FMR MV-FMR U-FMR
1
0.005 (0.007)
0.004 (0.005)
0.237 (0.465)
0.178 (0.282)
0.034 (0.009)
0.034 (0.010)
0.266 (0.006)
0.266 (0.006)
2
0.002 (0.001)
0.000 (0.001)
0.003 (0.001)
0.190 (0.293)
0.020 (0.138)
0.001 (0.010)
0.028 (0.050)
0.010 (0.018)
3
0.004 (0.005)
0.003 (0.002)
0.252 (0.405)
0.081 (0.181)
0.043 (0.012)
0.043 (0.012)
0.246 (0.006)
0.245 (0.006)
Table 1: Performance comparison of MV-FMR and U-FMR across simulation Scenarios 1-3, for continuous and binary outcomes. Values are presented as mean (SD) of the ISE for $\beta_{1}(t)$ (first line) and for $\beta_{2}(t)$ (second line).
We then evaluated the robustness of MV-FMR to different functional forms of the exposure-outcome relationships for Scenarios 1 and 3. Specifically, we considered three distinct combinations: (i) both exposures with linear effects, (ii) one exposure with a linear effect and the other with a quadratic effect, and (iii) both exposures with quadratic effects. Figure 3 presents the estimated coefficient functions $\beta_{1}(t)$ and $\beta_{2}(t)$ averaged across simulations for a continuous outcome, with corresponding results for a binary outcome provided in Supplementary Figure 1. Across all time points, MV-FMR consistently achieved lower mean squared error (MSE) than the univariable approach and more accurately recovered the true functional forms. Notably, MV-FMR successfully captured nonlinear time-varying effects, whereas the univariable failed to reproduce the underlying functional relationships. Additional simulations examined alternative exposure–outcome functional forms, including stepwise and decreasing lifetime effects (Supplementary Table 3). Together, these results confirm the robustness of MV-FMR across a wider range of true effect shapes.
Figure 3: Mean time-varying causal effect estimates at six time points for comparing MV-FMR and U-FMR with a continuous outcome. Columns represent different exposure-outcome model combinations, rows show the estimated values for $\beta_{1}(t)$ and $\beta_{2}(t$). Black crosses indicate the true causal effects, green stars show MV-FMR estimates, and orange triangles show U-FMR estimates. Top panel: Scenario 1. Bottom panel: Scenario 3.
We further investigated the performance of MV-FMR to different degrees of instrument overlap and causal structure. The results for the continuous outcome are presented in Table 2 and for the binary outcome in Supplementary Table 4. In Scenario 1, the proportion of overlapping instruments ($p_{12}$) was varied across 0.15, 0.3, 0.5, and 0.8 to evaluate performance under increasing pleiotropy. The ISE for both $\beta_{1}(t)$ and $\beta_{2}(t)$ remained consistently low across all levels of instrument overlap. In Scenario 3, we assessed sensitivity to mediation strength by varying the parameter $\gamma$ across values 0.15, 0.3, 0.5, and 0.8, which controls the strength of the causal pathway from $X_{1}(t)$ to $X_{2}(t)$. MV-FMR consistently recovered the direct causal effects with comparable ISE across all mediation strengths. These findings demonstrate that the method reliably distinguishes direct from mediated effects, even when the mediating pathway is strong.
Scenario 1
$p_{12}$ 0.15 0.30 0.50 0.80
0.005 (0.007)
0.003 (0.004)
0.004 (0.005)
0.004 (0.006)
0.004 (0.008)
0.004 (0.005)
0.004 (0.005)
0.005 (0.010)
Scenario 3
$\gamma$ 0.15 0.30 0.50 0.80
0.004 (0.005)
0.005 (0.006)
0.003 (0.003)
0.002 (0.002)
0.004 (0.005)
0.006 (0.007)
0.003 (0.002)
0.002 (0.002)
Table 2: MV-FMR performance under varying proportions of overlapping variants (Scenario 1) and different mediation strengths (Scenario 3) for continuous outcome. Values are the mean (SD) of the ISE for estimated $\beta_{1}(t)$ (first line) and $\beta_{2}(t)$ (second line).
Finally, we performed a sensitivity analysis to evaluate the effect of longitudinal data sparsity on MV-FMR performance. The number of observation points per individual exposure was varied across $\{5,10,25,50\}$ over the interval $[0,50]$. This analysis was conducted for Scenarios 1 and 3 with a continuous outcome, with corresponding results for a binary outcome provided in Supplementary Table 5. As shown in Table 3, MV-FMR performance improved with increasing data density, confirming that denser longitudinal sampling enhances the reliable recovery of functional relationships.
nSparse
5 10 25 50
Scenario 1
0.027 (0.037)
0.023 (0.030)
0.005 (0.007)
0.004 (0.005)
0.001 (0.001)
0.001 (0.001)
0.001 (0.000)
0.001 (0.000)
Scenario 3
0.031 (0.036)
0.016 (0.018)
0.004 (0.005)
0.003 (0.002)
0.001 (0.001)
0.001 (0.001)
0.001 (0.000)
0.001 (0.000)
Table 3: MV-FMR performance under varying longitudinal data sparsity for continuous outcome. Values are mean (SD) of the ISE for estimated $\beta_{1}(t)$ (first line) and $\beta_{2}(t)$ (second line) at different numbers of observation points per individual (nSparse) in Scenarios 1 and 3.
4 Case study: Systolic Blood Pressure and Body Mass Index as time-varying exposure on Coronary Artery Disease
We applied the proposed MV-FMR framework to investigate the time-varying causal effects of systolic blood pressure (SBP) and body mass index (BMI) on coronary artery disease (CAD) risk. This application provides a good case study for our functional extension of multivariable MR for several reasons. First, SBP and BMI are well-established cardiovascular risk factors with substantial phenotypic correlation [hall2015obesity, collaboration2016body], creating a scenario in which multivariable approaches offer advantages over univariable methods. Second, biological evidence indicates that BMI influences cardiovascular outcomes both directly and indirectly through blood pressure [seravalle2014obesity], establishing a mediation pathway that traditional univariable MR cannot disentangle. Indeed, cross-sectional multivariable MR studies have shown that approximately 27% of the BMI-CAD association operates through SBP [Gill2021], providing a well-characterized benchmark for validation. Third, both exposures exhibit time-varying effects over the lifecourse, with evidence suggesting that critical exposure windows for cardiovascular risk may differ between midlife and later life [Gnatiuc2017, duncan2021differential]. By jointly modeling SBP and BMI trajectories while accounting for their correlation and potential mediation, MV-FMR enables estimation of the direct time-varying causal effect of each exposure on CAD risk, allowing us to recover established causal patterns and uncover critical windows of effect that cannot be identified using standard cross-sectional or univariable approaches.
To address this question, we used data from UK Biobank, a large-scale prospective cohort study that recruited over 500,000 participants aged 40-69 years between 2006 and 2010 across the United Kingdom [Sudlow2015]. For this analysis, we restricted to participants who met the following criteria: (i) European ancestry, (ii) availability of genetic data, and (iii) at least two longitudinal measurements of both SBP and BMI between ages 50 and 70 years. After applying these criteria, 59,907 participants were included in the final sample. SBP and BMI trajectories were reconstructed from all available measurements in primary care records and UK Biobank baseline assessment visits within the 50-70 year window. Incident CAD was defined using first-occurrence ICD-10 codes I21-I25, covering acute myocardial infarction and chronic ischemic heart disease.
We investigated the effects of SBP and BMI on CAD using 274 and 199 genetic variants, respectively. Genetic instruments were obtained from large-scale genome-wide association study (GWAS) meta-analyses: for SBP were taken from Evangelou et al. [Evangelou2018], and for BMI from Winkler et al. [Winkler2015]. Five variants were associated with both exposures, representing overlapping instruments that may introduce pleiotropy between SBP and BMI. We then applied MV-FMR to jointly estimate the time-varying causal effects of SBP and BMI on CAD risk over ages 50-70 years. Individual exposure trajectories were modeled using FPCA with PACE, with the number of components chosen via our data-driven k-fold cross-validation procedure. Two principal components were selected for both exposures. For SBP, the first two components explained 77.0% and 16.5% of the variability, respectively, while for BMI, they explained 90.5% and 8.1%. The large proportion of explained variance indicates that trajectories of both exposures were well captured by the two-component FPCA representation.
Instrument strength was assessed using F-statistics for each principal component. For SBP, the first and second components yielded conditional F-statistics of 9.77 and 3.44, respectively, while for BMI, the corresponding values were 9.03 and 2.57. Although components fell below the conventional threshold for strong instruments ($F<10$), our simulation results indicate that functional MR approaches remain unbiased in such settings. To assess the benefits of joint modeling, we compared MV-FMR with U-FMR applied separately to each exposure. This comparison allowed us to evaluate whether simultaneous modeling of SBP and BMI revealed different causal effect patterns compared to the univariable analyses. Figure 4 presents the MV-FMR estimates. The analysis revealed distinct time-varying causal effects of SBP and BMI on CAD risk. Both exposures showed positive associations with CAD risk between ages 50 and 60, but their effects attenuated thereafter, becoming non-significant (95% bootstrap CI) beyond age 60. These findings suggest that the critical exposure window for SBP and BMI in relation to CAD risk is concentrated in the sixth decade of life.
The comparison between MV-FMR and U-FMR provides evidence consistent with established mediation patterns. For SBP, the multivariable and univariable analyses yielded broadly consistent results, both identifying similar effect magnitudes and time windows of significance. For BMI, MV-FMR estimates showed attenuation in the point estimates compared to the univariable approach, suggesting that part of BMI’s apparent effect on CAD operates indirectly through blood pressure pathways. Inspection of the confidence intervals reveals partial overlap between MV-FMR and U-FMR estimates for BMI. Notably, this overlapping is also observed in the traditional MV-MR analysis [Gill2021], where the univariable BMI effect (OR 1.49, 95% CI 1.39 to 1.60) and the multivariable effect (OR 1.34, 95% CI 1.24 to 1.45) showed overlapping confidence intervals.
Figure 4: Estimated time-varying causal effects of SBP and BMI on CAD risk from MV-FMR (top) and U-FMR (bottom). Blue curves show SBP effects with 95% bootstrap CI; red curves show BMI effects with 95% bootstrap CI.
5 Discussion
In this work, we introduced MV-FMR, a novel framework for functional Mendelian Randomization that enables the estimation of time-varying causal effects of multiple longitudinal exposures. While Mendelian Randomization has been widely used to strengthen causal inference in epidemiology, most applications assume constant effects over the life course and focus on a single exposure. Recent methodological advances have extended MR to either multivariable settings or functional exposures; to our knowledge, MV-FMR is the first approach that integrates both dimensions simultaneously. This extension allows researchers to investigate multiple related exposures, mediation pathways, and horizontal pleiotropy while preserving the continuous nature of exposure trajectories.
MV-FMR addresses important methodological gaps in the growing literature on longitudinal MR. Previous approaches have considered functional methods for single exposures [cao2016timevarying, Tian2024] or multivariable methods with discretized time [sanderson2022timevarying]. In contrast, our framework jointly models multiple exposure processes as continuous functions of time. It leverages FPCA to represent longitudinal exposures, accommodating sparse and irregular measurements, and implements a data-driven cross-validation procedure for selecting the optimal number of components. By explicitly accounting for overlapping genetic instruments, MV-FMR allows direct and indirect effects to be disentangled even in the presence of pleiotropy. The framework provides estimation procedures for both continuous and binary outcomes, with theoretical justification for identifiability under standard multivariable MR assumptions.
Extensive simulation studies demonstrated the robustness and accuracy of MV-FMR across a wide range of scenarios, including nonlinear effects, horizontal pleiotropy, and mediation. The method reliably recovered the true functional forms of causal effects and consistently outperformed univariable functional MR approaches. The use of a robust estimation procedure ensured that weak instruments do not bias causal effect estimates. We further conducted sensitivity analyses varying instrument overlap, mediation strength, exposure–outcome functional forms, and data sparsity, demonstrating the method’s robustness under diverse practical conditions.
The practical relevance of the proposed MV-FMR framework was demonstrated through an application to UK Biobank data, examining the time-varying causal effects of SBP and BMI on CAD risk. This analysis was intended primarily as a proof-of-concept, demonstrating how the proposed framework can recover well-established causal structures while extending them to a functional setting. The results highlighted ages 50–60 as a period during which both SBP and BMI exerted their strongest estimated effects, with attenuated effects at older ages. The comparison between multivariable and univariable functional MR provides evidence consistent with established mediation mechanisms. While SBP estimates were quite unchanged between univariable and multivariable models, BMI effects were attenuated in the joint analysis, indicating that part of BMI’s association with CAD is captured through blood pressure pathways. This pattern mirrors findings from traditional cross-sectional MV-MR, where partial mediation of the BMI-CAD relationship by SBP has been reported and where univariable and multivariable estimates similarly exhibit overlapping confidence intervals [Gill2021]. Taken together, these findings indicate that MV-FMR behaves as expected in the presence of mediation, while providing a functional representation of effects across age.
Several limitations should be acknowledged. MV-FMR requires longitudinal measurements for each exposure within the same age range, which may not be available in all datasets. The computational burden increases with the number of exposures, as functional modeling is more demanding than cross-sectional analysis. Additionally, measurement error in exposures or any violations of the core MR assumptions could bias the causal estimates. While our simulation studies indicate that MV-FMR performs well even with moderately weak instruments, increasing the complexity of the underlying time-varying effects can reduce instrument strength. As the number of functional components used to represent exposures grows, the effective conditional F-statistics can decrease, potentially introducing bias if they fall below a critical threshold. Accordingly, tailored simulation analyses are essential to evaluate whether the available instruments and data structure are sufficient to support reliable inference in a given application. Such pre-analysis assessment allows researchers to evaluate model feasibility and anticipate potential limitations in estimating causal effects. Finally, although our empirical application demonstrates the method’s feasibility, it was primarily intended as a proof of concept; broader applications are necessary to fully establish practical utility across diverse epidemiological contexts, and the choice and timing of outcome measurement relative to exposures should be carefully examined through extensive sensitivity analyses.
In summary, MV-FMR represents a methodological advance in functional Mendelian Randomization, offering a flexible, robust, and interpretable framework for estimating time-varying causal effects in multivariable settings. By integrating functional data analysis with multivariable MR, the method enables the identification of life-course critical periods, the disentangling of mediation pathways, and an improved understanding of the dynamic interplay between correlated risk factors. Future work should extend the framework to accommodate higher-dimensional exposures and investigate computational strategies for large-scale biobank applications. MV-FMR holds promise for generating richer causal insights and informing more precise preventive interventions in population health.
Data sharing
UKB data are accessible for health-related research in the public interest. This study used data from the UKB Resource under Application Number 102297.
Acknowledgements
The present research has been supported by MUR, grant Dipartimento di Eccellenza 2023- 2027. F. Ieva acknowledges the National Plan for NRRP Complementary Investments "Advanced Technologies for Human-centred Medicine" (PNC0000003).